A contractor rented a sander for 5 h and a polisher for 5h for $50. For another job, she rented the sander for 4 h and the polisher for 8 h. The rental fees for that job totalled $56. Find the hourly rate charged for each tool.
Answer by mananth(16946)

x= cost per hour sander
y= cost per hour polisher
Total sold
5 x + 5 y = 50 .............1

4 x + 8 y = 56 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 5
-40 x -40 y = -400
20 x + 40 y = 280
Add the two equations
-20 x = -120
/ -20
x = 6
plug value of x in (1)
5 x + 5 y = 50
30 + 5 y = 50
5 y = 50 -30
5 y = 20
y = 4
$6 cost per hour sander
$4 cost per hour polisher
m.ananth@hotmail.ca